OpenStudy (anonymous):

lim x→infinity of [ (sqrt of 4x^6 − x) / (x^3 + 9)]

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \frac{ \sqrt{4x^6-x} }{ x^3+9 }\]